OpenStudy (anonymous):

the perimeter of a rectangle is 146 m. The length is 1 m more than twice the width. Find the dimensions

OpenStudy (saifoo.khan):

Suppose the width be x. So the length be: 1+2x. Now, substitue them into the formula. Perimeter = 2(Length + Width) 146 = 2( (1+2x) + x)) Now solve for x to get the width.

OpenStudy (anonymous):

how do i solve for x

OpenStudy (saifoo.khan):

146 = 2( (1+2x) + x))

OpenStudy (saifoo.khan):

146 = 2(2+3x) Now.

OpenStudy (across):

This is how I work problems out: "the perimeter of a rectangle..." I read perimeter, thus it must have something to do with:\[P=2(l+w),\]where l and w are the length and width, respectively. "... is 146 m..." Then I know my formula should be modified to:\[146=2(l+w).\] "... The length is 1 m more than twice the width...." That said, I know that\[l=2w+1.\] "... Find the dimensions." Hmm, I think I need to solve the following system:\[146=2(l+w)\]\[l=2w+1\]Can you do that?